# YourMT3+: Multi-instrument Music Transcription with Enhanced Transformer Architectures and Cross-dataset Stem Augmentation

### v2

@@ 3 Model
In the YourMT3+ taxonomy, YMT3 models match MT3’s [4] architecture and training. YPTF+Single models use PerceiverTF (PTF) encoder with MT3’s single-channel decoder and stem augmentation. Our empirical finding demonstrates that PTF’s hierarchical attention with instrument-group sub-task queries enhances multi-instrument AMT in complex mixtures. YPTF.MoE replaces the encoder’s FFN with mixture of experts (MoE), enabling task-specific encodings in multi-dataset training. These models efficiently process MIDI tokens instead of piano-roll. Our multi-channel decoder assigns instrument groups per channel and masks loss for unannotated instruments, allowing training with incomplete labels. The final YPTF.MoE+Multi model integrates all these features.
@@ 3.3 Output Tokens
The center panel of Figure 1 shows the output sequence $Y^{\prime}$ with a maximum $N$ time steps, and the tokens representing MIDI-like events are listed in Section F of Supplemental Document. We followed the note sequence structure proposed in MT3 [4], and two modifications were made to the MT3 tokens: (a) unused velocity tokens except 0 and 1 were removed, and (b) programs 100 and 101 were reserved for singing voice (melody) and singing voice (chorus), respectively.
Compared to traditional piano-roll [12, 18, 10, 9, 5], the use of MIDI-like tokens [4] offers several advantages. First, it is more memory-efficient, using tokens for note onset, shift, and offset rather than hundreds of frames for silence as in piano-rolls. This efficiency also simplifies handling of multi-instrument data; piano-rolls require separate matrices for each instrument group, while the token approach needs only a program token for each change. Second, using tokens allows expansion of program vocabulary without a substantial memory growth, unlike piano-rolls. Lastly, tokens explicitly represent note onsets and offsets, avoiding the extra post-processing required for piano-rolls.
@@ 3.4 Decoder
In our implementation, the PTF encoder’s output hidden states are grouped by allocating two latents per channel. With group-linear projection, $k=26$ latents result in $k^{\prime}=13$ projected channels. The Multi decoder then independently decodes each of the $k^{\prime}$ inputs, producing $k^{\prime}$ sequences for each program using parallel decoders with shared parameters. We set the maximum sequence length to $N_{\text{single}}=1024$ (as in MT3 [4]) and $N_{\text{multi}}=256$. Potential truncation loss is discussed further in Supplemental B.6.
@@ 4.2 Cross-dataset Stem Augmentation
Stopping criteria In Line 5 of Algorithm 1, three criteria are established to stop the iterative mixing among stems. The first criterion is an exponential decay $S(j)$ that serves as the survival function defined as $S(j)=e^{-\tau j},$ where $\tau$ controls the surviving curve with respect to $j$-th iteration. The second criterion restricts $\hat{X}_{\text{ex}}$ to a length $L$, measured as sequence length post-tokenisation. The last criterion, $j>J$ with $J=5$ allows mixing up to 5 segments per base segment.
@@ 5.2 Results and Discussion
Performance on Pop Music: As seen at the bottom of Table 3, our model demonstrated competitive performance on the synthetic dataset [29] compared to other multi-AMT models. In Figure 3, our final model achieved 50 to 90% performance for most instruments, except for a few non-mainstream ones like chromatic percussion (c. perc) and synth pad (s.pad) in the synthetic dataset. However, a significant limitation emerged in its performance on commercial pop music recordings, as shown in Figure 3. Particularly for non-main instruments (excluding piano, bass, vocals, and drums), our models performed below 10%. This suggests potential biases introduced by training primarily on synthetic datasets, which may not fully cover the diverse timbres of pop music. Furthermore, except for the piano, all the pitched instruments showed a significant gap in the chroma-level metric, suggesting substantial octave errors and hinting that more varied pitch-shifting could be beneficial.

### v3

@@ 3 Model
In the YourMT3+ taxonomy, YMT3 models match MT3’s [4] architecture and training. YPTF+Single models use PerceiverTF (PTF) encoder with MT3’s single-channel decoder and stem augmentation (+). Our empirical finding demonstrates that PTF’s hierarchical attention with instrument-group sub-task queries enhances multi-instrument AMT in complex mixtures. YPTF.MoE replaces the encoder’s FFN with mixture of experts (MoE), enabling task-specific encodings in multi-dataset training. These models efficiently process MIDI tokens instead of piano-roll. Our multi-channel decoder assigns instrument groups per channel and masks loss for unannotated instruments, allowing training with incomplete labels. The final YPTF.MoE+Multi model integrates all these features.
@@ 3.3 Output Tokens
The center panel of Figure 1 shows the output sequence $Y^{\prime}$ with a maximum $N$ time steps, and the tokens representing MIDI-like events are listed in Supplemental F. As noted in Section 5.2, models trained with more fine-grained vocabulary consistently perform better. Therefore, we use MT3_FULL_PLUS for training and MT3_MIDI_PLUS only for comparison tests with previous work. Following the note sequence structure in MT3 [4], we made two modifications to the MT3 tokens: (a) unused velocity tokens, except 0 and 1, were removed, and (b) programs 100 and 101 were reserved for singing voice (melody) and singing voice (chorus), respectively.
Compared to traditional piano-rolls [12, 18, 10, 9, 5], MIDI-like tokens [4] offer several advantages: they are more memory-efficient by representing note onset, shift, and offset with tokens rather than hundreds of frames; they simplify multi-instrument data handling by expanding the program vocabulary without significant memory increase, whereas piano-rolls need large separate matrices for each instrument; and they explicitly represent linked note onsets and offsets, avoiding extra post-processing required for piano-rolls.
@@ 3.4 Decoder
In our implementation, the PTF encoder’s output hidden states are grouped by allocating two latents per channel—with group-linear projection, $k=26$ latents result in $k^{\prime}=13$ projected channels. The Multi decoder then independently decodes each of the $k^{\prime}$ inputs, producing $k^{\prime}$ sequences for each program using parallel decoders with shared parameters. We set the maximum sequence length to $N_{\text{single}}=1024$ (as in MT3 [4]) and $N_{\text{multi}}=256$. Potential truncation loss is discussed further in Supplemental B.6.
@@ 4.2 Cross-dataset Stem Augmentation
Stopping criteria In Line 4 of Algorithm 1, three criteria are established to stop the iterative mixing among stems. The first criterion is an exponential decay $S(j)$ that serves as the survival function defined as $S(j)=e^{-\tau j},$ where $\tau$ controls the surviving curve with respect to $j$-th iteration. The second criterion restricts $\hat{X}_{\text{ex}}$ to a length $L$, measured as sequence length post-tokenisation. The last criterion, $j>J$ with $J=5$ allows mixing up to 5 segments per base segment.
@@ 5.2 Results and Discussion
Performance on Pop Music: As seen at the bottom of Table 3, our model demonstrated competitive performance on the synthetic dataset [29] compared to other multi-AMT models. In Figure 3, our final model achieved 50 to over 90% performance for most instruments, except for a few non-mainstream ones like chromatic percussion (c. perc) and synth pad (s.pad) in the synthetic dataset. However, a significant limitation emerged in its performance on commercial pop music recordings, as shown in Figure 3. Particularly for non-main instruments (excluding piano, bass, vocals, and drums), our models performed below 10%. This suggests potential biases introduced by training primarily on synthetic datasets, which may not fully cover the diverse timbres of pop music. Furthermore, except for the piano, all the pitched instruments showed a significant gap in the chroma-level metric, suggesting substantial octave errors and hinting that more varied pitch-shifting could be beneficial.